THE AREA OF TRIANGLE FORMED BY THE TANGENT AT THE POINT (X,Y) TO THE CIRCLE WITH CENTER AT ORIGIN AND RADIUS K
(

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Equation of circle x2+y2=k2
let the point be (x1,y1)
Equation of line is x(x1)+y(y1)=k2
we know that area formed by the origin and line ax+by=c is c2/2ab
similarly area formed by the tangent=k4/2xy
Thanks and Regards
Shaik Aasif
askIITians faculty